# Quantum information scrambling in adiabatically-driven critical systems

### v1

@@ 3.3 Symmetry-breaking thermal states
Having systematically analyzed the mechanism of AQIS in the LMG model and its effectiveness in terms of the number of populated eigenstates, we turn now our attention to a different initial state. Moreover, we will study how the reported AQIS depends on the control parameter $g_{1}$, i.e. on whether the system is driven from the symmetry-breaking into the normal phase or not.
@@
where $Z$ is the partition function that ensures a proper normalization of the state, ${\rm Tr}[\hat{\rho}_{\beta,SB}]=1$, while $\left|\psi_{k,SB}(g_{0})\rangle\right.=\frac{1}{\sqrt{2}}(\left|\psi_{k,+}(g_{0})\rangle\right.+x_{k}\left|\psi_{k,-}(g_{0})\rangle\right.)$ being $x_{k}=\pm 1$ depending on the expectation value of $\hat{S}_{x}$. As for Eq. (12), we consider a maximally symmetry-breaking state with ${\rm Tr}[\hat{\rho}_{\beta,SB}\hat{S}_{x}]>0$ so that $P(S_{x})$ shows only non-zero probability for positive eigenvalues of $\hat{S}_{x}$ (cf. Fig. 6(a)), while $P(E)$ displays the standard exponentially decaying populations (cf. Fig. 6(b)). Hence, the state $\hat{\rho}_{\beta,SB}$ corresponds to a thermal state projected onto one of the two symmetry-breaking branches, i.e. to a generalized Gibbs state.
@@
We then analyze how the realization of the adiabatic cycle scrambles the initial information, depending on the value of $g_{1}$. As shown in Fig. 1, to ensure that the whole state is driven into the normal phase, $g_{1}>1$. For a final value $g_{1}<1$, only a fraction of excited states will enter the normal phase, thus rendering the AQIS ineffective as $\delta\varphi_{k}=0$ for those eigenstates that remain within the symmetry-breaking phase. The results are plotted in Fig. 6. As for the initial state given in Eq. (12), the final distribution $P(S_{x})$ largely differs from the initial one since both branches are populated, ensured by $g_{1}=1.25$. For other choices of $g_{1}$, the final $P(S_{x})$ may fail to be a balanced distribution. This is shown in Fig. 6(c) where the final expectation value $\langle\hat{S}_{x}(2\tau)\rangle={\rm Tr}[\hat{U}(2\tau)\hat{\rho}_{\beta,SB}\hat{U}^{\dagger}(2\tau)S_{x}]$ is plotted as a function of $\tau$ and for three different values of $g_{1}$. For $g_{1}=0.1$ the system remains in the symmetry-breaking phase during the whole cycle, and thus $\delta\varphi_{k}=0\ \forall k$ resulting in the absence of scrambling. For $g_{1}=0.65$, only a fraction of eigenstates are driven into the normal phase, partially scrambling the initial information. Finally, for $g_{1}=1.25$ the whole system enters in the normal phase, which corresponds to the AQIS discussed previously. To better visualize this effect, in Fig. 6(d) we show the average value of $\langle S_{x}(2\tau)\rangle$ in the range $\tau\in[10^{3},10^{4}]$ as a function of $g_{1}$, which behaves as an order parameter for the AQIS.
@@
Figure 7: AQIS for the QRM. (a) Initial energy probability distribution $P(E_{n})$ for $\left|\psi(0)\rangle\right.=\left|\alpha\rangle\right._{m}\left|0\rangle\right._{s}$ for $\Omega=100\omega$, initial rescaled coupling $g_{0}=2$ and $\alpha=5$. The normal phase finds itself above the critical excitation energy $E_{c}/\Omega=-1/2$. Upon the adiabatic protocol, the energy probability distribution is equivalent to the initial one. (b) Expectation value of the symmetry-breaking observable $\hat{x}$ after the adiabatic protocol of duration $\tau$ from $g_{0}=2$ to $g_{1}$, as indicated in the key, assuming the adiabatic approximation. (c) Average symmetry-breaking expectation value, in the range $\tau\in\{10^{3},10^{4}\}$, as a function of the final coupling $g_{1}$. For $g_{1}\lesssim 1.5$, the whole state enters in the normal phase, leading into a complete adiabatic information scrambling. Error bars correspond to the standard deviation for $\langle x(2\tau)\rangle$ as plotted in (b). See the main text for further details.
@@ 4 Quantum Rabi model
The QRM describes the coherent interaction between a spin-$1/2$ and a single bosonic mode and thus constitutes a fundamental model in the realm of light-matter interaction and of key relevance in quantum technologies. The Hamiltonian of the QRM can be written as
@@
where $\Omega$, $\omega$, and $\lambda$ correspond to the frequency of the qubit, bosonic mode, and interaction strength among them. Despite comprising just a single spin, this model has been shown to display ground-, excited-state, and even dynamical quantum phase transitions in the limit of $\Omega/\omega\rightarrow\infty$ [53, 51, 52, 54]. In this particular parameter limit, which plays the role of the standard thermodynamic limit where critical phenomena typically take place, the ground-state QPT occurs at $\lambda_{c}=\sqrt{\omega\Omega}/2$. Hence, for convenience, we introduce the rescaled interaction strength, $g=\lambda/\lambda_{c}$ so that $g_{c}=1$ as in the LMG.
The phase diagram of the critical QRM is divided into three regions. For $g<g_{c}$, a normal phase. For $g>g_{c}$, and energy $E_{k}(g)<E_{c}=-\Omega/2$, one finds the symmetry-breaking phase where the $\mathbb{Z}_{2}$ parity symmetry of the total number of excitations $\hat{\Pi}$ is spontaneously broken. Finally, for $g>g_{c}$ and $E_{k}(g)>E_{c}$, the symmetry is restored and the model is again in the normal phase. The excited-state QPT takes place at the critical energy $E_{c}=-\Omega/2$ [52]. Besides the microscopic details, the universal critical features are equivalent to those of the LMG model, with a similar energy spectrum as the one depicted in Fig. 1 (note however the difference in the location of the normal ($g<g_{c}$) and symmetry-breaking phases ($g>g_{c}$)).
In the following, we exemplify that the AQIS also applies to the QRM. For that, we choose
@@
as our initial symmetry-breaking state. There $\left|\alpha\rangle\right._{m}=\hat{D}(\alpha)\left|0\rangle\right._{m}=\exp\left[\alpha\hat{a}^{\dagger}-\alpha^{*}\hat{a}\right]\left|0\rangle\right._{m}$ represents a coherent state with amplitude $\alpha$ and $\left|0\rangle\right._{m}$ the vacuum state, while the spin is its ground state, $\hat{\sigma}_{z}\left|0\rangle\right._{s}=-\left|0\rangle\right._{s}$. This corresponds to a state in other of the wells of the effective double-well potential in the symmetry-breaking phase [51, 52]. The symmetry-breaking information is then attributed to the expectation value of the position operator $\hat{x}=(\hat{a}^{\dagger}+\hat{a})/\sqrt{2}$.
The adiabatic protocol is performed from $g_{0}=2$ to $g_{1}$, according to Eq. (5). Relying on the adiabatic approximation and choosing $\alpha=5$ for a critical QRM with $\Omega=100\omega$, we compute the final state and obtain $\left.\langle\psi(2\tau)\right|\hat{x}\left|\psi(2\tau)\rangle\right.$ as a function of $g_{1}$. Numerical simulations have been performed with $1000$ Fock states. The results are gathered in Fig. 7. First, the energy probability distribution is shown in Fig. 7(a), indicating that the state is contained within the symmetry-breaking phase since all the populated eigenstates have energies below the critical one $E_{c}=-\Omega/2$. As before, the AQIS becomes effective when the whole state is driven into the normal phase. For the considered initial state, this happens for $g_{1}\lesssim 1.5$ (cf. Fig. 7(b) and (c)). Note that this corresponds to a larger parameter than its critical value $g_{c}=1$ where the QPT takes place. Instead, the value $g\approx 1.5$ marks the position when the energy of the adiabatically-driven state surpasses the critical energy at which an excited-state QPT occurs [52].
@@ 5 Conclusions
In this article, we have analyzed the adiabatic scrambling of the symmetry-breaking information encoded in an initial state. The mechanism for such adiabatic quantum information scrambling is detailed, which is related to the difference of accumulated phases among degenerated eigenstates due to the driving from a symmetry-breaking into a normal phase, and thus to phase transitions taking place both in the ground and excited states of the system. Owing to its adiabatic nature, the protocol does not alter either the energy or the expectation values of observables independent of the symmetry-breaking information. To the contrary, as a consequence of the scrambling, the expectation value of an operator that quantifies the symmetry-breaking becomes approximately zero provided the initial state has a large support in the eigenstates of the Hamiltonian and the relative difference among the accumulated phases of eigenstates of opposite parity are uniformly distributed. We showcase the adiabatic quantum information scrambling in the Lipkin-Meshkov-Glick and quantum Rabi models and quantify the effectiveness of the scrambling for different initial states. The potential information retrieval is studied utilizing the Loschmidt echo, which indicates that an effective scrambling would require an almost perfect time-reversal protocol to restore the initial information. Our results demonstrate the intriguing interplay between quantum information, many-body systems, and critical phenomena.
Acknowledgments
F.J.G- R is grateful for financial support from the Spanish MCIN with funding from the European Union Next Generation EU (PRTRC17.I1) and the Consejería de Educación from Junta de Castilla y Leon through the QCAYLE project, as well as grants PID2020-113406GB-I00 MTM funded by AEI/10.13039/501100011033, and RED2022-134301-T

### v2

@@ 3.3 Symmetry-Breaking Thermal States
Having systematically analyzed the mechanism of AQIS in the LMG model and its effectiveness in terms of the number of populated eigenstates, we turn now our attention to a different initial state. Moreover, we will study how the reported AQIS depends on the control parameter $g_{1}$, i.e., on whether the system is driven from the symmetry-breaking into the normal phase or not.
@@
where $Z$ is the partition function that ensures a proper normalization of the state, ${\rm Tr}[\hat{\rho}_{\beta,SB}]=1$, while $\left|\psi_{k,SB}(g_{0})\rangle\right.=\frac{1}{\sqrt{2}}(\left|\psi_{k,+}(g_{0})\rangle\right.+x_{k}\left|\psi_{k,-}(g_{0})\rangle\right.)$ and $x_{k}=\pm 1$, depending on the expectation value of $\hat{S}_{x}$. As for Equation (12), we consider a maximally symmetry-breaking state with ${\rm Tr}[\hat{\rho}_{\beta,SB}\hat{S}_{x}]>0$ so that $P(S_{x})$ shows only non-zero probability for positive eigenvalues of $\hat{S}_{x}$ (cf. Figure 6a), while $P(E)$ displays the standard exponentially decaying populations (cf. Figure 6b). Hence, the state $\hat{\rho}_{\beta,SB}$ corresponds to a thermal state projected onto one of the two symmetry-breaking branches, i.e., to a generalized Gibbs state.
@@
We then analyze how the realization of the adiabatic cycle scrambles the initial information, depending on the value of $g_{1}$. As shown in Figure 1, to ensure that the whole state is driven into the normal phase, $g_{1}>1$. For a final value $g_{1}<1$, only a fraction of excited states will enter the normal phase, thus rendering AQIS ineffective, as $\delta\varphi_{k}=0$ for those eigenstates that remain within the symmetry-breaking phase. The results are plotted in Figure 6. As for the initial state given in Equation (12), the final distribution $P(S_{x})$ largely differs from the initial one, since both branches are populated, ensured by $g_{1}=1.25$. For other choices of $g_{1}$, the final $P(S_{x})$ may fail to be a balanced distribution. This is shown in Figure 6(c), where the final expectation value $\langle\hat{S}_{x}(2\tau)\rangle={\rm Tr}[\hat{U}(2\tau)\hat{\rho}_{\beta,SB}\hat{U}^{\dagger}(2\tau)S_{x}]$ is plotted as a function of $\tau$, and for three different values of $g_{1}$. For $g_{1}=0.1$ the system remains in the symmetry-breaking phase during the whole cycle, and thus $\delta\varphi_{k}=0\ \forall k$, resulting in the absence of scrambling. For $g_{1}=0.65$, only a fraction of eigenstates are driven into the normal phase, partially scrambling the initial information. Finally, for $g_{1}=1.25$, the whole system enters in the normal phase, which corresponds to AQIS, as discussed previously. To better visualize this effect, in Figure 6(d), we show the average value of $\langle S_{x}(2\tau)\rangle$ in the range $\tau\in[10^{3},10^{4}]$ as a function of $g_{1}$, which behaves as an order parameter for AQIS.
@@
Figure 6: AQIS for a symmetry-breaking thermal state in the LMG model. (a) Initial probability distribution $P(S_{x})$ over the observable $\hat{S}_{x}$ for $\hat{\rho}_{\beta,SB}$ with $\beta=0.02$ for $N=10^{3}$ and $g_{0}=0$ (solid black line), together with the probability for the state upon an adiabatic protocol for $\tau=10^{3}$ and $g_{1}=1.25$ (blue squares), which equally populates both $S_{x}$ branches. (b) Population of the initial eigenstate for the symmetry-breaking thermal state $\hat{\rho}_{\beta,SB}$, as in (a). (c) Final expectation value of $\hat{S}_{x}$ after an adiabatic protocol from $g_{0}=0$ to $g_{1}=0.1$ (red), $0.65$ (green), and $1.25$ (blue), as a function of $\tau$ and for $\hat{\rho}_{\beta,SB}$ with $\beta=0.02$ and $N=10^{3}$, as in (a,b). For $g_{1}=0.1$, the system is not driven into the normal phase. For $g=0.65$, the state enters partially, while for $g_{1}=1.25$ it is driven completely into the normal phase, leading to a full AQIS. (d) Average expectation value of $\hat{S}_{x}$ in the range $\tau\in\{10^{3},10^{4}\}$ as a function of the final coupling strength $g_{1}$, revealing an order-parameter-like behavior. The error bars correspond to a standard deviation for the average of $\langle\hat{S}_{x}(2\tau)\rangle$. See the main text for details.
Figure 7: AQIS for the QRM. (a) Initial energy probability distribution $P(E_{n})$ for $\left|\psi(0)\rangle\right.=\left|\alpha\rangle\right._{m}\left|0\rangle\right._{s}$ for $\Omega=100\omega$, initial rescaled coupling $g_{0}=2$, and $\alpha=5$. The normal phase finds itself above the critical excitation energy $E_{c}/\Omega=-1/2$. Upon the adiabatic protocol, the energy probability distribution is equivalent to the initial one. (b) Expectation value of the symmetry-breaking observable $\hat{x}$ after the adiabatic protocol of duration $\tau$ from $g_{0}=2$ to $g_{1}$, as indicated in the key, assuming the adiabatic approximation. (c) Average symmetry-breaking expectation value, in the range $\tau\in\{10^{3},10^{4}\}$, as a function of the final coupling $g_{1}$. For $g_{1}\lesssim 1.5$, the whole state enters in the normal phase, leading into a complete adiabatic information scrambling. Error bars correspond to the standard deviation for $\langle x(2\tau)\rangle$, as plotted in (b). See the main text for further details.
@@ 4 Quantum Rabi Model
The QRM describes the coherent interaction between a spin-$1/2$ and a single bosonic mode, and thus constitutes a fundamental model in the realm of light-matter interaction and of key relevance in quantum technologies. The Hamiltonian of the QRM can be written as
@@
where $\Omega$, $\omega$, and $\lambda$ correspond to the frequency of the qubit, bosonic mode, and interaction strength among them. Despite comprising just a single spin, this model has been shown to display ground-, excited-state, and even dynamical quantum phase transitions in the limit of $\Omega/\omega\rightarrow\infty$ [65, 63, 64, 66]. In this particular parameter limit, which plays the role of the standard thermodynamic limit where critical phenomena typically take place, the ground-state QPT occurs at $\lambda_{c}=\sqrt{\omega\Omega}/2$. Hence, for convenience, we introduce the rescaled interaction strength, $g=\lambda/\lambda_{c}$, so that $g_{c}=1$, as in the LMG.
The phase diagram of the critical QRM is divided into three regions. For $g<g_{c}$, one has a normal phase. For $g>g_{c}$, and energy $E_{k}(g)<E_{c}=-\Omega/2$, one finds the symmetry-breaking phase where the $\mathbb{Z}_{2}$ parity symmetry of the total number of excitations $\hat{\Pi}$ is spontaneously broken. Finally, for $g>g_{c}$ and $E_{k}(g)>E_{c}$, the symmetry is restored, and the model is again in the normal phase. The excited-state QPT takes place at the critical energy $E_{c}=-\Omega/2$ [64]. Besides the microscopic details, the universal critical features are equivalent to those of the LMG model, with a similar energy spectrum as the one depicted in Figure 1 (note, however, the difference in the location of the normal ($g<g_{c}$) and symmetry-breaking phases ($g>g_{c}$)).
In the following, we exemplify that AQIS also applies to the QRM. For that, we choose
@@
as our initial symmetry-breaking state. There, $\left|\alpha\rangle\right._{m}=\hat{D}(\alpha)\left|0\rangle\right._{m}=\exp\left[\alpha\hat{a}^{\dagger}-\alpha^{*}\hat{a}\right]\left|0\rangle\right._{m}$ represents a coherent state with amplitude $\alpha$, and $\left|0\rangle\right._{m}$ is the vacuum state, while the spin is its ground state, $\hat{\sigma}_{z}\left|0\rangle\right._{s}=-\left|0\rangle\right._{s}$. This corresponds to a state in the other well of the effective double-well potential in the symmetry-breaking phase [63, 64]. The symmetry-breaking information is then attributed to the expectation value of the position operator, $\hat{x}=(\hat{a}^{\dagger}+\hat{a})/\sqrt{2}$.
The adiabatic protocol is performed from $g_{0}=2$ to $g_{1}$, according to Equation (5). Relying on the adiabatic approximation, and choosing $\alpha=5$ for a critical QRM with $\Omega=100\omega$, we compute the final state, and obtain $\left.\langle\psi(2\tau)\right|\hat{x}\left|\psi(2\tau)\rangle\right.$ as a function of $g_{1}$. Numerical simulations have been performed with $1000$ Fock states. The results are gathered in Figure 7. First, the energy probability distribution is shown in Figure 7a, indicating that the state is contained within the symmetry-breaking phase, since all of the populated eigenstates have energies below the critical one, $E_{c}=-\Omega/2$. As before, AQIS becomes effective when the whole state is driven into the normal phase. For the considered initial state, this happens for $g_{1}\lesssim 1.5$ (cf. Figure 7b,c). Note that this corresponds to a larger parameter than its critical value $g_{c}=1$, where the QPT takes place. Instead, the value $g\approx 1.5$ marks the position when the energy of the adiabatically driven state surpasses the critical energy at which an excited-state QPT occurs [64].
@@ 5 Conclusions
In this article, we have analyzed the adiabatic scrambling of the symmetry-breaking information encoded in an initial state. The mechanism for such adiabatic quantum information scrambling is detailed, which is related to the difference in accumulated phases among degenerated eigenstates due to the driving from a symmetry-breaking phase into a normal phase, and thus to phase transitions taking place both in the ground and excited states of the system. Owing to its adiabatic nature, the protocol does not alter either the energy or the expectation values of observables independent of the symmetry-breaking information. To the contrary, as a consequence of the scrambling, the expectation value of an operator that quantifies the symmetry-breaking becomes approximately zero, provided that the initial state has large support in the eigenstates of the Hamiltonian, and the relative difference among the accumulated phases of eigenstates of opposite parity are uniformly distributed. We showcase the adiabatic quantum information scrambling in the Lipkin–Meshkov–Glick and quantum Rabi models, and quantify the effectiveness of the scrambling for different initial states. The potential information retrieval is studied utilizing the Loschmidt echo, which indicates that an effective scrambling would require an almost perfect time-reversal protocol to restore the initial information. Our results demonstrate the intriguing interplay between quantum information, many-body systems, and critical phenomena.
Author Contributions:
Conceptualization, R.P. and F.J.G.-R.; Formal analysis, R.P. and F.J.G.-R.; Investigation, R.P. and F.J.G.-R.; Writing – original draft, R.P. and F.J.G.-R. All authors have read and agreed to the published version of the manuscript.
Funding:
F.J.G- R is grateful for financial support from the Spanish MCIN with funding from the European Union Next Generation EU (PRTRC17.I1) and the Consejería de Educación from Junta de Castilla y Leon through the QCAYLE project, as well as grants PID2023-148409NB-I00 MTM funded by AEI/10.13039/501100011033, and RED2022-134301-T.
Data Availability Statement:
The raw data supporting the conclusions of this article will be made available by the authors on request.
Conflicts of Interest:
The authors declare no conflicts of interest.